6 - 5[8-(2y-4)]
can you help me? I am not sure if the brackets are multiplied or not.
I came up with -54-10y and 8-(2y-4)

Found 2 solutions by tutorcecilia, stanbon:
Answer by tutorcecilia(2152)

I agree with your first answer.
6 - 5[8-(2y-4)]
6 - 5[8-2y+4]
6 - 5[12-2y]
6 - 60+10y
-54+10y

Answer by stanbon(75887)

6 - 5[8-(2y-4)]
Work from the inner-most bracket outwards, as follows:
=6-5[8-2y+4]
=6-5[12-2y]
=6-60+10y
=-54+10y
Cheers,
Stan H.